OpenStudy (anonymous):

WILL GIVE MEDAL The table below represents the amount of money Cathy has in her money box as a function of time Time (month) x Money ($) y 0 400 1 320 2 240 3 160

OpenStudy (anonymous):

Part A: What is the y-intercept of the function and what does this tell you about the table? (4 points) Part B: What is the average rate of change of the function represented by the table between x = 1 to x = 3 months and what does the average rate represent? (4 points) Part C: What is the domain of the function represented by the table if Cathy does not put any money into the box but takes out the same amount of money from

OpenStudy (anonymous):

Time (month) x Money ($) y 0 400 1 320 2 240 3 160

OpenStudy (anonymous):

@HelpBlahBlahBlah @raeror @johnweldon1993

OpenStudy (anonymous):

@helpout3 @Emily_13

OpenStudy (anonymous):

PLEASE HELP WILL GIVE MEDALS

OpenStudy (anonymous):

@Laine6789 @ranga

OpenStudy (anonymous):

@JaxsonH22

OpenStudy (anonymous):

@cordell10

OpenStudy (anonymous):

@DVReed96 @taylorlee @mathslover @mathmale

OpenStudy (anonymous):

Yo mention more people i can help need all 3 answers?

OpenStudy (anonymous):

yes please thank you so much :D

OpenStudy (anonymous):

I'm sorry I'm new I'm lost

OpenStudy (anonymous):

it's alright i just need the answers and the fastest way to get them thank you for looking tho

OpenStudy (anonymous):

Idk how to leave

OpenStudy (anonymous):

click a different question to leave. @DVReed96 can you help/ do you know how?

OpenStudy (anonymous):

im working on it

OpenStudy (anonymous):

No? I can help but Idk how on this

OpenStudy (anonymous):

need help asap

OpenStudy (anonymous):

this is difficult to do on a computer man

OpenStudy (anonymous):

ik that is why i posted it XD

OpenStudy (anonymous):

Y-intercept = 400

OpenStudy (anonymous):

I need to find out a graphing program otherwise i cant do much

OpenStudy (anonymous):

I believe you r supposed to be able to just use the table

OpenStudy (anonymous):

@EmeraldSunset ?

OpenStudy (anonymous):

@Brittney_kay12 @FibonacciChick666 @taffalousswagg_bru

OpenStudy (anonymous):

OK...what was the question again?

OpenStudy (anonymous):

Part A: What is the y-intercept of the function and what does this tell you about the table? (4 points) Part B: What is the average rate of change of the function represented by the table between x = 1 to x = 3 months and what does the average rate represent? (4 points) Part C: What is the domain of the function represented by the table if Cathy does not put any money into the box but takes out the same amount of money from

OpenStudy (anonymous):

I would need the table..

OpenStudy (anonymous):

ok, nm you already attached it

OpenStudy (anonymous):

XD

OpenStudy (anonymous):

Are there any options? Or is this a free response question?

OpenStudy (anonymous):

Free response :,C

OpenStudy (anonymous):

@taylorlee

OpenStudy (anonymous):

ok, figured out some of it

OpenStudy (anonymous):

first you have to find the slope

OpenStudy (anonymous):

slope (m) = y2 - y1/x2 - x1

OpenStudy (anonymous):

m = 2 - 1/1 - 0

OpenStudy (anonymous):

m = 1/1 m = 1

OpenStudy (anonymous):

Part A: What is the y-intercept of the function and what does this tell you about the table? (4 points) Part B: What is the average rate of change of the function represented by the table between x = 1 to x = 3 months and what does the average rate represent? (4 points) Part C: What is the domain of the function represented by the table if Cathy does not put any money into the box but takes out the same amount of money from

OpenStudy (anonymous):

sorry, that wasn't right bear with me, I haven't done this kind of prob. in a while

OpenStudy (anonymous):

So the yintercet and slope are 1?

OpenStudy (anonymous):

y - y1 = m(x - x1) y - 2 = m(x - 1) y - 2 = 1(x - 1) y - 2 = x - 1 y (-2 + 2) = x (-1 + 2) y = x + 1 Find the y-intercept by substituting x with 0 y = 0 + 1 y = 1

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

y=400-80x domain is {0,1,2,3}

OpenStudy (anonymous):

gods, that took forever

OpenStudy (johnweldon1993):

@Tman9876 y-intercept = where on the graph x = 0 So at x = 0...we have y = 400 right?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I really need a direct answer not around the bush stuff just tell me and show me how u did it please (show ur work) thanks

OpenStudy (anonymous):

Part A: What is the y-intercept of the function and what does this tell you about the table? (4 points) Part B: What is the average rate of change of the function represented by the table between x = 1 to x = 3 months and what does the average rate represent? (4 points) Part C: What is the domain of the function represented by the table if Cathy does not put any money into the box but takes out the same amount of money from

OpenStudy (anonymous):

IDK

OpenStudy (elise_a18):

dang this is hard :I sorry bro, my head is about to explode

OpenStudy (anonymous):

its ok everyone keeps saying that

OpenStudy (anonymous):

what school do you go to anyways?

OpenStudy (anonymous):

(A) Y-Intercept = 1 (copy and paste all the work I did into the box, you don't want to get points off for not showing work) The y intercept represents where x = 0

OpenStudy (anonymous):

Moving on to B

OpenStudy (anonymous):

rate of change is the same as the slope, so again your answer would be 1

OpenStudy (anonymous):

Average rate of change is the speed at which a variable changes over a specific period of time

OpenStudy (anonymous):

Moving on to C

OpenStudy (anonymous):

C doesn't look complete. Would you mind posting the entire question? :)

OpenStudy (fibonaccichick666):

@tman9876 Did you get it?